# A geometric perspective on Algebraic Quantum Field Theory

### v1

@@ 3.3 Regularity and Localizability
Given a connected Lie group $G$, such that $\mathcal{E}({\mathfrak{g}})\neq\{0\}$, a ${\mathbb{Z}}_{2}$-graded extension $G_{\tau_{h}}$ with respect to an Euler involution $\tau_{h}$, where $h\in\mathcal{E}({\mathfrak{g}})$, and an (anti-)unitary representation $U$ of $G_{\tau_{h}}$, one can construct a net of standard subspaces on wedges as in Sect. 2.4.3. The Bisognano–Wichmann property holds, which naturally raises the question of the extent to which the regularity property (b) from Theorem 31 is also satisfied. This is a representation theoretic question investigated in [MN24] that in part uses the results of [FNÓ23].
In this setting we reformulate the property as follows. Assuming that $h$ is an Euler element and $(U,\mathcal{H})$ an (anti-)unitary representation $(U,\mathcal{H})$ of $G_{\tau_{h}}$. Let ${\tt V}$ be the standard subspace associated to the wedge $(h,\tau_{h})$ by the BGL–construction. We shall say that $U$ is regular with respect to $h$, or $h$-regular, if there exists an $e$-neighborhood $N\subseteq G$ such that ${\tt V}_{N}=\bigcap_{g\in N}U(g){\tt V}$ is cyclic. Replacing $N$ by its interior, we may always assume that $N$ is open.
@@
We now present the state of the art on the analysis of this property:
Regularity via positive energy
In conformal Algebraic Quantum Field Theory, the geometry of the model ensures that one can shrink wedge regions: conformal symmetries in Minkowski space can transform a wedge region $W$ into a doublecone region $O\Subset W$ and both regions are associated 1-1 to Euler elements; for Möbius symmetries of the circle, intervals are associated to Euler wedges, and interval inclusions are associated to inclusions of Euler wedges [BGL93, BGL02, MN21]. In particular, provided the BGL construction for these models, there exists an open neighborhood of the identity $N\subset G$ such that given a $\bigcap_{g\in N}gW_{1}$ contains a wedge region supporting a cyclic subspace, hence regularity for the BGL net holds.
Firstly, we situate the general framework within the Lie theory context, where the regularity property stems from the existence of a generating positive cone and the positivity of the energy. Theorem [MN24, Thm. 4.9] claims that, if $(U,\mathcal{H})$ is an (anti-)unitary representation of $G_{\tau_{h}}$ for which the cones
$C_{\pm}=\pm C_{U}\cap{\mathfrak{g}}_{\pm 1}(h)$
@@
As examples one can consider the Möbius group. Here, the condition on the cone $C_{\pm}$ to be generating holds for positive energy representations of the Möbius group. Up to sign, the only pointed, generating closed convex $\mathop{{\rm Ad}}\nolimits$-invariant cone has been introduced in (5). Given a positive energy representation $U$ of Möb, then $C_{U}=C$. Considering the Euler element ${h=\frac{1}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}}$, we have
$C_{\pm}=\pm C\cap{\mathfrak{g}}_{\pm 1}(h),\quad C_{+}={\mathbb{R}}_{+}\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\quad C_{-}={\mathbb{R}}_{+}\begin{pmatrix}0&0\\ 1&0\end{pmatrix},$
@@
Another example is the the following. Let ${\mathfrak{g}}=\mathop{{\mathfrak{so}}}\nolimits(2,d+1)$ the Lie algebra of the conformal group of the Minkowski spacetime ${\mathbb{R}}^{1,d}$, consider the Euler element $h\in{\mathfrak{g}}$ generating the dilations $\exp(ht)x=e^{t}x$ where $x\in{\mathbb{R}}^{1,d}$, then ${\mathfrak{g}}_{1}\simeq{\mathbb{R}}^{1,d}$ and the positive cone $C$ is generating in ${\mathfrak{g}}_{1}$ (and analogously for ${\mathfrak{g}}_{-1}$). The generating property holds for all hermitian simple real Lie algebras, see [NÓØ21, Lem. 3.2].
Regularity in a semidirect product
As a second step, it looks natural to study the regularity property for (anti-)unitary representation of semidirect products. This picture now contains the Poincaré group that is not simple. In general, a simply connected Lie group $G$, is a semidirect product $G\cong R\rtimes S$, where $S$ is semisimple and $R$ is the solvable radical (Levi decomposition, see e.g. [Ho81, Thm. VIII.4.3]).
One can split the problem checking two conditions on the two subgroups defining the semidirect product. Consider a semidirect product Lie group $G=R\rtimes L$, let $0\neq h\in\mathcal{E}({\mathfrak{g}})$, and $(U,\mathcal{H})$ be an (anti-)unitary representation of $G_{\tau_{h}}$. Theorem 4.11 in [MN24] proves that $(U,\mathcal{H})$ is regular under the following two conditions
@@
    $(U|_{L},\mathcal{H})$ is regular.
This theorem applies to (anti-)unitary positive energy representation of the Poincaré group $\mathcal{P}_{+}$ as follows: look at the representation of the Poincaré group $\mathcal{P}_{+}^{\uparrow}={\mathbb{R}}^{1,d}\rtimes\mathcal{L}_{+}^{\uparrow}$, check the non-triviality of the one-dimensional cones $C_{\pm}$ in the eigenspaces ${\mathfrak{r}}_{\pm 1}(h_{W_{R}})$ that corresponds to the lightrays ${\mathbb{R}}({\bf{e}}_{0}\pm{\bf{e}}_{1}){\mathfrak{r}}\cong{\mathbb{R}}^{1,d}$ and the regularity property for the restriction of the representation to the identity component $\mathcal{L}_{+}^{\uparrow}$ of the Lorentz group. The first property corresponds to the positive energy condition on the Poincaré representation. The second one holds for every representation of the Lorentz group, as one can see from Theorem 4.25 in [MN24]. Note that one can relate this special case of the results in the Appendix of [BGL02, Thm 4.7]
To study the regularity property for reductive Lie groups $G$ one can introduce the localizability property. Given an (anti-)unitary representation $(U,\mathcal{H})$ of the Lie group $G$, this property ensures the existence of a concrete cyclic subspace of the representation Hilbert space described in terms of distribution vectors of the Hilbert space $\mathcal{H}$, localized on an open region of a causal symmetric manifold $M$ lying in the intersection of wedge subspaces (wedge regions as described in Sect. 2.3). One explicitly constructs local subspaces (in analogy with the the free field case) and proves the cyclicity of subspaces associated to intersections of wedge regions and the Bisognano–Wichmann property for standard subspace of wedge regions of the causal manifold. This has been studied in [FNÓ23] and recalled in this setting to conclude regularity in [MN24]. Results in [MN24, Thm. 4.23, Cor 4.24] guarantee localizability for representations of a connected reductive Lie group. Then the above theorem applies to the cases where $G=R\rtimes L$, $U|_{R}$ satisfies (a) and $L$ is reductive.
Currently, it is unknown whether all (anti-)unitary representations of Lie groups of the form $G_{\tau_{h}}$, where $h\in\mathcal{E}({\mathfrak{g}})$ is an Euler element, are regular. The preceding discussion suggests that resolving this question requires a more detailed analysis of the case of solvable groups.
@@ 3.4 The type of wedge algebras
Von Neumann algebras can be classified in terms of the geometry of their projection and the spectrum of the modular action [Ta02, Su87, Co73]. Algebras taking part in an AQFT are proved to be type III1 in many cases and it is expected to be always the case when they are nontrivial (see for instance [Dr77, Lo82, Fr85, BDF87, BB99, FG94]). Recently, type II1 algebras have become increasingly relevant in certain constructions within the AQFT context, see [CLPW23, FJLRW25].
We prove that our generalized AQFT setting also respects this expectation. Let $G$ be a connected Lie group with Lie algebra ${\mathfrak{g}}$. We say that $h$ is anti-elliptic if ${\mathfrak{n}}_{h}+{\mathbb{R}}h={\mathfrak{g}}$ where ${\mathfrak{n}}_{h}\trianglelefteq{\mathfrak{g}}$ be the smallest ideal of ${\mathfrak{g}}$ such that the image of $h$ in the quotient Lie algebra ${\mathfrak{g}}/{\mathfrak{n}}_{h}$ is elliptic, namely $\mathop{{\rm ad}}\nolimits x$ is semisimple with purely imaginary spectrum. Let $(U,\mathcal{H})$ be a unitary representation of $G$ with discrete kernel, $\mathcal{N}\subset\mathcal{M}\subseteq B(\mathcal{H})$ an inclusion of von Neumann algebras, and $\Omega\in\mathcal{H}$ a unit vector which is cyclic and separating for $\mathcal{N}$ and $\mathcal{M}$. Assume that the following properties hold:
@@
If, in addition, $\mathcal{H}^{G}={\mathbb{C}}\Omega\not=\mathcal{H}$, then $\mathcal{M}$ is factor of type III1. This theorem is [MN24, Thm. 5.15].
Note that if $h\in\mathcal{E}({\mathfrak{g}})$ is anti-elliptic, then ${\mathfrak{g}}_{0}\subseteq{\mathbb{R}}h+[{\mathfrak{g}}_{1},{\mathfrak{g}}_{-1}]$. Anti-elliptic elements are referred to as essential element in [St08]. One can apply this setting when it is considered a unitary representation $U$ on an Hilbert space $\mathcal{H}$ of the connected Lie group $G$ with discrete kernel, a vector $\Omega\in\mathcal{H}$ that is the unique unit vector fixed by $U(G)$ and $\mathcal{M}\subseteq B(\mathcal{H})$ be a von Neumann algebra for which $\Omega$ is cyclic and generating. Assume that $h\in{\mathfrak{g}}$ is anti-elliptic and the following properties.
@@
    Modularity: There exists an element $h\in{\mathfrak{g}}$ for which $e^{2\pi i\partial U(h)}=\Delta_{\mathcal{M},\Omega}$. As $\ker(U)$ is discrete, $h$ is uniquely determined.
@@
Then $h$ is an Euler element and $\mathcal{M}$ is a type III1 factor. This picture includes the cases of nets of von Neumann algebras on Euler wedges when, under the previous assumptions,
$\mathcal{W}_{+}\ni W\mapsto\mathcal{M}(gW)=U(g)\mathcal{M}(W_{h})U(g)^{-1}\subset\mathcal{B}(\mathcal{H}),\qquad\mathcal{M}(W_{h})=\mathcal{M},\,W_{h}=(h,\tau_{h}),\,g\in G$
defines a $G$-covariant isotonous net of von Neumann algebras.
@@
We have a measurable decomposition $X=X_{0}\dot{\cup}X_{1}$, where $\mathop{{\rm dim}}\nolimits\mathcal{H}_{x}=1$ for $x\in X_{0}$ and the representations $(U_{x})_{x\in X_{0}}$ are trivial. For $x\in X_{1}$, the algebras $\mathcal{M}_{x}$ are factors of type III1 and $(\mathcal{M}_{x},\Omega_{x},\underline{U}_{x})$ satisfies (Reg) and (Mod), where $\underline{U}_{x}$ is the representation of $G/\ker(U_{x})$ induced by $U_{x}$, further details in [MN24, Thm. 5.22].
@@ 4 An outlook on hermitian Lie algebra
We have seen that various aspects of the geometric framework of Algebraic Quantum Field Theory (AQFT) can be explored at the Lie theory level. In particular, the Bisognano–Wichmann (BW) property and the modular reflection lead to a wedge-boost generalizing correspondence when a Lie group symmetry provided with an Euler element of the Lie algebra is considered. As we have seen along these lines, this framework allows for the deduction of properties of wedge regions, wedge symmetries, wedge standard subspaces and wedge von Neumann algebras.
Within this generalized setting, conformal models stand out due to the significant role played by orthogonal wedges, offering a richer structure [MN21]. Here the Lie algebra of the symmetry group is hermitian, namely a Lie algebra such that the center of a maximal compactly embedded subalgebra ${\mathfrak{k}}\subset{\mathfrak{g}}$ is non-zero. A simple real Lie algebra ${\mathfrak{g}}$ is hermitian if and only if it has a convex $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$-invariant cone $C\neq\{0\}$ or ${\mathfrak{g}}$. Furthermore a simple real Lie algebra containing Euler elements is hermitian if and only if it is of tube type, i.e. the restricted root system is of type $C_{r}$ (see [MN21, Prop.3.11]). In view of the classification in Sect.2.2, in this case, there exists a unique orbit of symmetric Euler elements. In [MN21, Theorem 3.13] in a simple Lie algebra symmetric Euler elements have orthogonal Euler element partners. On Minkowksi space this is depicted by wedges in orthogonal directions, as $W_{i}=\{x\in{\mathbb{R}}^{1,d},|x_{0}|<x_{i}\}$ with $i=1,2$, where $j_{W_{1}}\Lambda_{W_{2}}(t)j_{W_{1}}=\Lambda_{W_{2}}^{-1}(t)$. In our language orthogonal Euler elements are the Lie elements defining orthogonal wedges, satisfy $\tau_{h_{1}}(h_{2})=-h_{2}$, where $h_{1},h_{2}\in\mathcal{E}({\mathfrak{g}})$.
One can consider couples of orthogonal elements and study orbits of orthogonal elements. The existence of disjoint orbits is particularly relevant in Conformal Field Theory on Minkowksi spacetime. Given a unitary representation $U$ of the universal covering of the conformal group $\widetilde{G}$, where ${\mathfrak{g}}=\mathop{{\mathfrak{so}}}\nolimits(2,d+1)$, inequivalent couples of orthogonal Euler elements generate inequivalent representation of $\widetilde{\mathop{{\rm SL}}}\nolimits_{2}({\mathbb{R}})$ not all of them for instance have positivity of the energy. A systematic analysis of these orthogonal pairs, along with a geometric examination of the conformal models, will be discussed in forthcoming papers together with an analysis of their relationship with conformal AQFT models ([MNO25, MN25]).

### v2

@@ 3.3 Regularity and Localizability
Given a connected Lie group $G$, such that $h\in\mathcal{E}({\mathfrak{g}})$, and an (anti-)unitary representation $U$ of $G_{\tau_{h}}$, one can construct a net of standard subspaces on wedges as in Sect. 2.4.3. The Bisognano–Wichmann property holds, which naturally raises the question of the extent to which the regularity property (b) from Theorem 31 is also satisfied. This is a representation theoretic question investigated in [MN24] that in part uses the results of [FNÓ23].
In this setting we reformulate the property as follows. Assuming that $h$ is an Euler element and $(U,\mathcal{H})$ an (anti-)unitary representation of $G_{\tau_{h}}$. Let ${\tt V}$ be the standard subspace associated to the wedge $(h,\tau_{h})$ by the BGL–construction. We shall say that $U$ is regular with respect to $h$, or $h$-regular, if there exists an $e$-neighborhood $N\subseteq G$ such that ${\tt V}_{N}=\bigcap_{g\in N}U(g){\tt V}$ is cyclic. Replacing $N$ by its interior, we may always assume that $N$ is open.
@@
We now present the state of the art on the analysis of this property.
3.3.1 Regularity via generating positive cone
In conformal Algebraic Quantum Field Theory, the geometry of the model ensures that one can shrink wedge regions: conformal symmetries in Minkowski space can transform a wedge region $W$ into a doublecone region $O\Subset W$ and both regions are associated 1-1 to Euler elements; for the Möbius wedge geometry, intervals are associated to Euler wedges, and interval inclusions are associated to inclusions of Euler wedges [BGL93, BGL02, MN21]. In particular, provided the BGL construction for these models, there exists an open neighborhood of the identity $N\subset G$ such that given a $\bigcap_{g\in N}gW_{1}$ contains a wedge region supporting a cyclic subspace, hence regularity for the BGL net holds.
Firstly, we situate the general framework within the Lie theory context, where the regularity property stems from the existence of a generating positive cone. Theorem [MN24, Thm. 4.9] claims that, if $(U,\mathcal{H})$ is an (anti-)unitary representation of $G_{\tau_{h}}$ for which the cones
$C_{\pm}:=\pm C_{U}\cap{\mathfrak{g}}_{\pm 1}(h)$ (31)
@@ 3.3.1 Regularity via generating positive cone
As examples one can consider the Möbius group. Here, the condition (31) holds for positive energy representations of the Möbius group. Up to sign, the only pointed, generating closed convex $\mathop{{\rm Ad}}\nolimits$-invariant cone has been introduced in (7). Given a positive energy representation $U$ of Möb, then $C_{U}=C$. Considering the Euler element ${h=\frac{1}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}}$, we have
$C_{\pm}:=\pm C\cap{\mathfrak{g}}_{\pm 1}(h),\quad C_{+}={\mathbb{R}}_{+}\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\quad C_{-}={\mathbb{R}}_{+}\begin{pmatrix}0&0\\ 1&0\end{pmatrix},$
@@
Another example is the the following. Let ${\mathfrak{g}}=\mathop{{\mathfrak{so}}}\nolimits(2,d+1)$ the Lie algebra of the conformal group of the Minkowski spacetime ${\mathbb{R}}^{1,d}$, consider the Euler element $h\in{\mathfrak{g}}$ generating the dilations $\exp(ht)x=e^{t}x$ where $x\in{\mathbb{R}}^{1,d}$, then ${\mathfrak{g}}_{1}\simeq{\mathbb{R}}^{1,d}$ and the positive cone $C_{+}=\{x\in{\mathbb{R}}^{1,d}:x^{2}\geq 0,x_{0}\geq 0\}$ is generating in ${\mathfrak{g}}_{1}$ (and analogously for ${\mathfrak{g}}_{-1}$). The generating property holds for all hermitian simple real Lie algebras, see [NÓØ21, Lem. 3.2].
3.3.2 Regularity in a semidirect product
As a second step, it looks natural to study the regularity property for (anti-)unitary representations of semidirect products where one can split the problem checking two conditions on the two defining subgroups. This picture now contains the Poincaré group that is not simple. Furthermore any simply connected Lie group $G$, is a semidirect product $G\cong N\rtimes S$, where $S$ is semisimple and $N$ is the solvable radical (Levi decomposition, see e.g. [HN12, Thm. 11.1.19]).
We consider a general (connected) semidirect product Lie group $G=R\rtimes L$, let $h\in\mathcal{E}({\mathfrak{g}})$, and $(U,\mathcal{H})$ be an (anti-)unitary representation of $G_{\tau_{h}}$. Theorem 4.11 in [MN24] proves that $(U,\mathcal{H})$ is regular under the following two conditions
@@ 3.3.2 Regularity in a semidirect product
    $(U|_{L_{\tau_{h}}},\mathcal{H})$ is $h$-regular.
This theorem applies to (anti-)unitary positive energy representations of the Poincaré group $G_{\tau_{W_{R}}}$. In this context let $R={\mathbb{R}}^{1,d}$, $L=\mathcal{L}_{+}^{\uparrow}$ and $h=h_{W_{R}}\in\mathcal{E}({\mathfrak{g}})$, one can check the generating property of the one-dimensional cones $C_{\pm}$ in the eigenspaces ${\mathfrak{r}}_{\pm 1}(h_{W_{R}})$ that corresponds to the lightrays ${\mathbb{R}}({\bf{e}}_{0}\pm{\bf{e}}_{1}){\mathfrak{r}}={\mathfrak{r}}_{\pm 1}$ and the regularity property for the restriction of the representation to the identity component $\mathcal{L}_{+}^{\uparrow}$ of the Lorentz group. The first property follows from the spectral condition, namely positivity of the energy of the the Poincaré representation. The second one holds for every representation of the Lorentz group, as one can see from Theorem 4.25 in [MN24].
To study the regularity property for reductive Lie groups $G$ one can introduce the localizability property. Given an (anti-)unitary representation $(U,\mathcal{H})$ of the Lie group $G_{\tau_{h}}$, this property ensures the existence of a concrete cyclic real subspace of the representation Hilbert space localized on an open region of a causal symmetric manifold $M$ lying in the intersection of wedge subspaces (wedge regions as described in Sect. 2.3). In [FNÓ23], the authors explicitly constructs local subspaces (in analogy with the the free field case) and prove the Bisognano–Wichmann property for standard subspaces of wedge regions and the cyclicity of the subspaces associated to open regions that are intersections of wedge regions. This result is recalled to conclude regularity in [MN24]. Results in [MN24, Thm. 4.23, Cor 4.24] guarantee localizability for representations of a connected reductive Lie group. Then Theorem 4.11 in [MN24] applies to the cases where $G=R\rtimes L$, $U|_{R}$ satisfies (a) and $L$ is reductive.
Currently, it is unknown whether all (anti-)unitary representations of Lie groups of the form $G_{\tau_{h}}$ are regular. The preceding discussion suggests that resolving this question requires a more detailed analysis of the case of solvable groups.
@@ 3.4 The type of wedge algebras
Von Neumann algebras can be classified in terms of the geometry of their projection and the spectrum of the modular action [Ta02, Su87, Co73]. Algebras involved in an AQFT are proven to be of type III1 in many cases, and they are expected to always be of this type when nontrivial (see for instance [Dr77, Lo82, Fr85, BDF87, BB99, FG94]). Recently, type II1 algebras have become increasingly relevant in certain constructions within the AQFT context, see [CLPW23, FJLRW25].
We prove that our generalized AQFT setting also respects this type III1 expectation. Let $G$ be a connected Lie group with Lie algebra ${\mathfrak{g}}$. We say that $h$ is anti-elliptic if ${\mathfrak{n}}_{h}+{\mathbb{R}}h={\mathfrak{g}}$ where ${\mathfrak{n}}_{h}\trianglelefteq{\mathfrak{g}}$ be the smallest ideal of ${\mathfrak{g}}$ such that the image of $h$ in the quotient Lie algebra ${\mathfrak{g}}/{\mathfrak{n}}_{h}$ is elliptic. Let $(U,\mathcal{H})$ be a unitary representation of $G$ with discrete kernel, $\mathcal{N}\subset\mathcal{M}\subseteq B(\mathcal{H})$ an inclusion of von Neumann algebras, and $\Omega\in\mathcal{H}$ a unit vector which is cyclic and separating for $\mathcal{N}$ and $\mathcal{M}$. Assume that the following properties hold:
@@
If, in addition, $\mathcal{H}^{G}={\mathbb{C}}\Omega\not=\mathcal{H}$, then $\mathcal{M}$ is a factor (i.e. $\mathcal{M}\cap\mathcal{M}^{\prime}={\mathbb{C}}\textbf{1}$) of type III1. This theorem is [MN24, Thm. 5.15].
Note that if $h\in\mathcal{E}({\mathfrak{g}})$ is anti-elliptic, then ${\mathfrak{g}}_{0}\subseteq{\mathbb{R}}h+[{\mathfrak{g}}_{1},{\mathfrak{g}}_{-1}]$. Anti-elliptic elements are called essential elements in [St08]. One can apply this setting when it is considered a unitary representation $U$ on an Hilbert space $\mathcal{H}$ of the connected Lie group $G$ with discrete kernel, a vector $\Omega\in\mathcal{H}$ that is the unique unit vector fixed by $U(G)$ and $\mathcal{M}\subseteq B(\mathcal{H})$ be a von Neumann algebra for which $\Omega$ is cyclic and separating. Assume that $h\in{\mathfrak{g}}$ is anti-elliptic and the following properties.
@@
    Modularity: There exists an element $h\in{\mathfrak{g}}$ for which $e^{2\pi i\partial U(h)}=\Delta_{\mathcal{M},\Omega}$.
@@
Then $h$ is an Euler element and $\mathcal{M}$ is a type III1 factor. This picture includes the cases of nets of von Neumann algebras on Euler wedges when, under the previous assumptions, the map
$\mathcal{W}_{+}\ni W\mapsto\mathcal{M}(gW)=U(g)\mathcal{M}(W_{h})U(g)^{-1}\subset\mathcal{B}(\mathcal{H}),\qquad g\in G$
where $\mathcal{M}(W_{h})=\mathcal{M}$ and $\,W_{h}=(h,\tau_{h}),$ defines a $G$-covariant isotonous net of von Neumann algebras.
@@
We have a measurable decomposition $X=X_{0}\dot{\cup}X_{1}$, where $\mathop{{\rm dim}}\nolimits\mathcal{H}_{x}=1$ for $x\in X_{0}$ and the representations $(U_{x})_{x\in X_{0}}$ are trivial. For $x\in X_{1}$, the algebras $\mathcal{M}_{x}$ are factors of type III1 and $(\mathcal{M}_{x},\Omega_{x},\underline{U}_{x})$ satisfies (Reg) and (Mod), where $\underline{U}_{x}$ is the representation of $G/\ker(U_{x})$ induced by $U_{x}$. Further details in [MN24, Thm. 5.22].
@@ 4 An outlook on hermitian Lie algebra
We have seen that various aspects of the geometric framework of Algebraic Quantum Field Theory (AQFT) can be explored at the Lie theory level. In particular, the Bisognano–Wichmann and modular reflection properties give rise to a generalized wedge-boost correspondence, where the Euler elements assume a central role. As we have seen along these lines, this framework allows for the deduction of properties of wedge regions, wedge symmetries, wedge standard subspaces and wedge von Neumann algebras.
Within this generalized setting, conformal models stand out due to the significant role played by orthogonal wedges, offering a rich structure [MN21]. Here the Lie algebra of the symmetry group is assumed to be simple and hermitian. Specifically a simple Lie algebra is hermitian if the center of a maximal compactly embedded subalgebra ${\mathfrak{k}}\subset{\mathfrak{g}}$ is non-zero. Equivalently, a simple real Lie algebra ${\mathfrak{g}}$ is hermitian if and only if it has a convex $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$-invariant cone $C\neq\{0\}$ or ${\mathfrak{g}}$. Furthermore a simple hermitian real Lie algebra contains an Euler element if and only if it is of tube type, i.e. the restricted root system is of type $C_{r}$ (see [MN21, Prop.3.11]). In view of the classification in Sect.2.2, in the hermitian simple case there exists a unique $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$-orbit of symmetric Euler elements. By [MN21, Theorem 3.13], in a simple Lie algebra symmetric Euler elements have orthogonal Euler element partners: We shall say that $h_{a},h_{b}\in\mathcal{E}({\mathfrak{g}})$ in a Lie algebra ${\mathfrak{g}}$ are orthogonal Euler elements if they satisfy the condition
$\tau_{h_{a}}(h_{b})=-h_{b}.$
As examples, this is the case of the Euler elements associated to the right half-circle and the upper half-circle of the wedge geometry on $S^{1}$ and the case of Euler elements associated to ${\mathbb{R}}^{1,d}$-Minkowksi wedges $W_{i}=\{x\in{\mathbb{R}}^{1,d},|x_{0}|<x_{i}\}$ with $i=1,2$ and $d\geq 2$.
One can consider couples of orthogonal elements in a simple Hermitian Lie algebra and study orbits of orthogonal elements. The existence of disjoint orbits is particularly evident in Conformal Field Theory on Minkowksi spacetime. Given a unitary representation $U$ of the universal covering of the conformal group $\widetilde{G}$, where ${\mathfrak{g}}=\mathop{{\mathfrak{so}}}\nolimits(2,d+1)$, inequivalent couples of orthogonal Euler elements generate inequivalent representations of $\widetilde{\mathop{{\rm SL}}}\nolimits_{2}({\mathbb{R}})$ not all of them satisfying a non-trivial spectral condition. A systematic analysis of these orthogonal pairs, along with a geometric examination of the conformal models, will be discussed in forthcoming papers together with an analysis of their relationship with conformal AQFT models ([MNO25, MN25]).